CLICK HERE TO REVIEW LEARNING OBJECTIVES 2 Not complete Marked out of 5.00 F Flag question Compute, Disaggregate and Interpret ROE and ROA Selected balance sheet and income statement information from Staples, Inc., follows ($ millions). Sales Interest expense Net income Total Assets Stockholders Equity 2014 $22,492 2014 2014 2014 2013 2014 2013 $49 $135 $10,314 $11,175 $5,313 $6,141 a. Compute the 2014 return on equity (ROE), return on assets (ROA), and return on financial leverage (ROFL) Round answers to one decimal place (ie.. 0.2568-25.7%). Do not round until your final answer. ROE ROA ROFL b. Disaggregate ROA into profit margin (PM) and asset turnover (AT Round answers to one decimal place (i.e., 0.2568-25.7%). Do not round until your final answer. PM AT Check
Answer #1

Please provide rating..

Answer a)
ROE = net income/Average equity
ROE = =135/((5313+6141)/2)
ROE = 2.4%
ROA = Net income/Average Asset
ROA= =135/((10314+11175)/2)
ROA= 1.3%
ROFL = Net income before interest and taxes/(Average Debt)
we have to compute the averate debt and net income before interest and taxes
Net income before interest and tax = Net income + interest expenses
=135+149
284 mil
** its assumed that we don thave any tax
Computation of average debt
Year 2013 2014
i Total asset 11175 10314
ii Total equity 6141 5313
iii=i-ii Debt 5034 5001
Average debt = =(5034+5001)/2 5017.50 mil
ROFL = =284/5017.5 5.7%
Answer b)
PM = Profit margin = Net income/ Total sales =
=135/22492
0.6%
AT = Asset turnover = Total sales/ Average asset
=22492/((10314+11175)/2)
         2.09
Prof = We have computed above ROA = 1.3%
If we multiply PM and AT ratio we will get ROA
See below -
PM*AT = 0.6%*2.09 1.3% Which is exactly same as ROA